# Dynamic Tuning Towards Parameter and Inference Efficiency for ViT Adaptation

### v1

@@ Abstract
Existing parameter-efficient fine-tuning (PEFT) methods have achieved significant success on vision transformers (ViTs) adaptation by improving parameter efficiency. However, the exploration of enhancing inference efficiency during adaptation remains underexplored. This limits the broader application of pre-trained ViT models, especially when the model is computationally extensive. In this paper, we propose Dynamic Tuning (DyT), a novel approach to improve both parameter and inference efficiency for ViT adaptation. Specifically, besides using the lightweight adapter modules, we propose a token dispatcher to distinguish informative tokens from less important ones, allowing the latter to dynamically skip the original block, thereby reducing the redundant computation during inference. Additionally, we explore multiple design variants to find the best practice of DyT. Finally, inspired by the mixture-of-experts (MoE) mechanism, we introduce an enhanced adapter to further boost the adaptation performance. We validate DyT across various tasks, including image/video recognition and semantic segmentation. For instance, DyT achieves comparable or even superior performance compared to existing PEFT methods while evoking only $71\%-85\%$ of their FLOPs on the VTAB-1K benchmark.
Keywords:
Vision Transformers (ViTs) Parameter Efficient Fine-tuning (PEFT) Dynamic Tuning (DyT)
@@ 1 Introduction
With the remarkable success of Vision Transformers (ViTs) [19, 48, 25], fine-tuing pre-trained ViT on other data domains [75] or task applications [74, 34, 54, 66] is becoming a common practice. However, as model sizes increase [73, 15, 47], the associated adaptation costs become prohibitive due to the burden of fine-tuning and inference on the target task. Parameter-efficient fine-tuning (PEFT) methods (e.g. AdaptFormer [10], LoRA [32], and VPT [34]), are proposed to tackle the tuning problem by reducing tunable model parameters. They usually update a small amount of parameters while keeping the original model fixed, which reduces learnable parameters effectively while maintaining fine-tuning accuracy.
@@
Figure 1: FLOPs and Accuracy of ViT-B/16 [19] on VTAB-1K [74]. “Full tuning” denotes that all parameters are trainable during fine-tuning. AdaptFormer [10], LoRA [32] and VPT [34] are typical PEFT methods.
Dynamic networks [69, 58, 44] demonstrate that token pruning helps to reduce model inference costs. However, these designs primarily focus on pre-training from scratch or full tuning on the same dataset, without considering data domain transfer. Furthermore, the token pruning operation constraints these methods on visual recognition scenarios, limiting their further applications e.g. dense prediction [77]. Our motivation thus arises from how to develop dynamic modules together with PEFT methods to enhance both parameter and inference efficiency during ViT adaptation, which should also suit a wide range of visual tasks.
@@
Comprehensive evaluations for DyT are conducted from multiple perspectives. In the image domain, as shown in Figure 1(b). DyT surpasses existing PEFT methods while consuming only $85\%$ of the ViT-B FLOPs on the VTAB-1K benchmark [74]. When visual tokens are scaled up from images to videos, our DyT shows superior generalization ability on action recognition benchmarks, e.g. K400 [9] and SSV2 [24], with a reduction of 37GFLOPs. In the scenario where labels are scaled up from recognition to segmentation, our DyT even outperforms full tuning on ADE20K [77] with 21GFLOPs reduction. These experimental results indicate that the proposed DyT is efficient in both parameter and inference across various visual tasks and directs a promising field for efficient model adaptation.
@@ 2 Related Works
Parameter-efficient fine-tuning (PEFT) is designed to adapt a pre-trained model to downstream tasks by only tuning a small part of the parameters. Existing PEFT methods can broadly be categorized into three groups: adapter-based methods, re-parametrization-based methods, and prompt-based methods. Adapter-based methods [10, 31, 56, 36, 12] insert some tiny modules into the original model and only these inserted modules are updated during fine-tuning. Re-parametrization approaches [72, 32, 43, 20, 17] usually cooperate with reparameterization techniques, directly modifying specific parameters in the pre-trained model. Prompt-based methods [34, 2, 79, 78, 38, 8, 7] involve appending a small number of learnable tokens to input sequences, thereby adapting intermediate features in frozen layers. Please refer to [70, 42] to find more comprehensive reviews.
@@
Dynamic neural networks[27] can dynamically adjust their architectures based on the input data, which enables them to control the computational redundancy based on input data. Existing methods have explored dynamic layer depth [61, 4, 68, 29, 26], dynamic channel width [30, 40, 28] and dynamic routing [41] in convolutional neural networks. When it comes to the era of vision transformer, many works [64, 60, 58, 44, 52] attempt to improve the inference efficiency by reducing the token redundancy. For instance, Liang et al. [44] identify and consolidate inattentive tokens into only one token in each transformer layer. Rao et al. [58] progressively discard less informative tokens between layers. Wang et al. [64] adjust the number of tokens to represent images. A more detailed literature review can be found in [27, 63]. Although these approaches have shown significant success in vision tasks, they require training a model from scratch or fine-tuning all parameters on the same dataset used in pre-training, making them unsuitable for efficient ViT adaptation.
@@ 3.1 Preliminary
The adapter architectures have been widely explored in previous PEFT methods [10, 36, 31]. An adapter is generally designed as a bottleneck module, which consists of a project layer $\mathbf{W}_{down}\in\mathbb{R}^{C\times d}$ to squeeze the channel dimension and another project layer $\mathbf{W}_{up}\in\mathbb{R}^{d\times C}$ for dimension restoration. Here, $d$ denotes the bottleneck dimension and $d\ll C$ ensures the efficiency of the adapter. Given an input feature $\mathbf{X}$, the adapter operation is formulated as:
@@ 3.2 Dynamic Tuning
In Vision Transformers (ViTs), the computational burden primarily resides in the transformer layers. Specifically, the operations in $\operatorname{Attn}$ and $\operatorname{MLP}$ account for approximately $35.8\%$ and $63.5\%$ of total FLOPs in ViT-B/16 [19]. Previous works [58, 44] have revealed that there exists a token-redundancy issue in ViTs and found that some tokens can be discarded without sacrificing the performance. Inspired by this, we propose an efficient ViT adaptation approach, named Dynamic Tuning (DyT), which not only maintains parameter efficiency during fine-tuning but also reduces redundant computation during inference. The core idea lies in dynamically selecting tokens for processing within transformer blocks. Given the input tokens $\mathbf{X}$ of a block, DyT can be formulated as:
@@
Figure 2: Overview of Dynamic Tuning. (a) In the fine-tuning stage, we adopt Gumbel Noise to enable end-to-end training. (b)In the inference stage, $\operatorname{TD}$ selects $K$ activated tokens $\mathbf{X}_{s}$ from $\mathbf{X}$ based on the mask $\mathbf{M}$, which saves the computations on those deactivated tokens in $\operatorname{Block}$. $\operatorname{Block}$ can represent a multi-head self-attention block $\operatorname{Attn}$, a multi-layer perceptron $\operatorname{MLP}$, or an entire transformer layer.
@@
However, directly using threshold makes $\mathbf{M}$ a discrete decision, resulting in a non-differentiable problem during fine-tuning. To address this, we incorporate the Gumbel Noise [30] into sigmoid to replace the original sigmoid function during fine-tuning. It can be formulated as:
@@
where $\mathbf{G}_{1}$, $\mathbf{G}_{2}\sim\operatorname{Gumbel}(0,1)$. $\tau$ represents the temperature and is set to 5.0 by default. Further details of Gumbel-Sigmoid are provided in the supplementary material. Subsequently, we can obtain $\mathbf{\overline{M}}$ using the same operation in Equation.3. The Gumbel Noise makes the sampling of $\mathbf{\overline{M}}$ stochastic during training and we adopt $\mathbf{\overline{S}}$ as a differentiable approximation of $\mathbf{\overline{M}}$. Both of these help $\operatorname{TD}$ to be trained in an end-to-end manner. The calculation of forward and backward propagations during training can be formulated as:
@@
Figure 3: Model variants. For brevity, we omit the LayerNorm [1] in $\operatorname{Attn}$ and $\operatorname{MLP}$ blocks. “DyT” denotes the dynamic tuning presented in Figure 2.
@@ 3.4 MoE-Adapter
In DyT, the adapter is responsible for processing all tokens, requiring it to have enough capability, particularly when the downstream tasks e.g. semantic segmentation are challenging to adapt. To tackle this problem, we propose a MoE-adapter, inspired by mixture-of-experts [59, 67]. It effectively enhances the capability of the adapter with introducing negligible computation cost.
@@ 3.5 Loss Functions
For an image $I$, we calculate the cross-entropy loss $\mathcal{L}_{cls}=-\mathbf{y}\operatorname{log}(\mathbf{\hat{y}})$ between the predicted probability $\mathbf{\hat{y}}$ and its corresponding one-hot label $\mathbf{y}$, to supervise the learning of classification. To control the average activation rate $r$ in the entire model, we add a loss term to constrain the number of activated tokens in the dynamic tuning, which can be formulated as:
$\mathcal{L}_{rate}=(\frac{1}{L\times N}\sum_{l=1}^{L}\sum_{n=1}^{N}\mathbf{M}_{n}^{l}-r)^{2},$ (8)
where $L$ denotes the number of layers in ViT. $\mathbf{M}_{n}^{l}$ represents the mask generated in $\operatorname{TD}$ from the $l$-th layer. Finally, the loss function is defined as $\mathcal{L}=\mathcal{L}_{cls}+\alpha\mathcal{L}_{rate}$, where $\alpha$ serves as the weight of the activation rate loss and is set to 2.0 by default.
@@ 4.1 Experiment Settings
To evaluate the adaptation performance, we conduct experiments on VTAB-1K [74] benchmark. The training data in this benchmark is extremely limited, with only 1,000 training samples for each task. We also conduct experiments on three image classification datasets with complete training sets, including CIFAR-100 [39], SVHN [23], Food-101 [5]. We adopt two video datasets, Kinetic-400 (K400) [9] and Something-Something V2 (SSv2) [24], to evaluate the performance when the number of tokens scaled up. For the dense prediction task, we evaluate our method on two widely recognized semantic segmentation datasets, AED20K [77] and COCO-stuff [6].
@@
We conduct all experiments based on the ViT-B/16 [19], which is pre-trained on ImageNet21K [16] with full supervision, unless otherwise specified. The bottleneck dimension $d$ is set to 64 by default. We adopt the same training schedule as [10]. Detailed hyperparameters for each experiment can be found in the supplementary material. The default setting in experiments is marked in color.
@@ 4.2 Analysis
In Table 1, we compare the performance of four model variants across both image and video datasets. We set the activation rate $r$ in “Attention Dispatch” and “MLP Dispatch” variants to 0.5, and to 0.7 for “Attention-MLP Dispatch” and “Layer Dispatch” variants, and train each respectively. This results in similar average FLOPs for four variants. We observe that the default setting “MLP Dispatch” achieves superior performance across four datasets while maintaining the lowest computational cost. Although “Attention-MLP Dispatch” and “Layer Dispatch” exhibit slightly better performance on K400, the former incurs double additional parameters while the latter lacks generalization capability on other datasets. The comparison between “MLP Dispatch” and other variants proves that only skipping tokens in MLP blocks is a better design.
@@
(a) CIFAR-100
(b) SVHN
(c) Food-101
Investigations of dispatch level and dispatch strategy.
Token Sample TD Random CIFAR-100 SVHN Food-101 K400 SSv2
91.37 97.08 90.32 74.39 45.34
90.70 97.29 89.8 71.14 44.09
89.38 96.79 86.39 68.27 40.13
87.15 97.07 85.26 68.18 39.79
Different bottleneck dimensions.
We investigate the impact of the bottleneck dimensions of the adapter in DyT and the results are summarized in Table 3, revealing that different datasets prefer different configurations. When the dataset is easy to adapt, such as CIFAR-100, a bottleneck dimension of $d=4$ is sufficient to achieve satisfactory performance. Conversely, video datasets require larger adapter dimensions e.g. $d=256$ to attain better performance. We set the default bottleneck dimension to 64, which achieves a balance between performance and cost, such as the number of additional parameters and FLOPs. It is worth noting that the relationship between FLOPs and bottleneck dimension is not strictly monotonic. For instance, when $d=16$, the FLOPs are lower than when $d=1$, since the training may converge at activating fewer tokens as the number of adapter’s parameters increases.
CIFAR-100 SVHN Food-101 K400 SSv2
1 0.03 12.11 91.46 95.5 89.65 71.78 36.22
4 0.08 11.89 91.57 96.61 89.99 73.16 39.24
16 0.30 12.06 91.46 96.98 90.24 73.65 42.49
64 1.19 12.21 91.37 97.08 90.32 74.39 45.34
256 4.73 13.32 91.18 97.03 90.42 75.33 46.73
@@
We observe that the performance on image datasets drops when we increase the expert number in MoE-adapter. This phenomenon can be attributed to the simplicity of image datasets and the model does not require too many parameters to adapt. In contrast, for video datasets, such as K400 and SSv2, the best accuracy is achieved with 4 and 8 experts, respectively. The reason behind this is that the domain gap between the pre-training dataset and video datasets is large and the model needs sufficient adapter capacity to learn the adaptation. This proves that we can introduce the MoE-adapter when the target dataset is difficult to adapt.
@@
In Figure 6, we visualize the token activation rates across different layers in ViT-B/16 [19]. We observe that the model tends to activate more tokens in lower layers while deactivating tokens in higher layers. This phenomenon can be attributed to that the model tends to take more general knowledge from the lower layers of the pre-trained model and learn more task-specific knowledge in higher levels. Additionally, the activation results vary across different datasets. For instance, SSv2 demonstrates increased token activation rate in Layer 5 and Layer 6 compared to other datasets, whereas SVHN experiences substantial token deactivation in Layer 6, 7, and 8. This discrepancy arises from that the model needs different knowledge from the pre-trained weights to address dataset-specific challenges.
It is noteworthy that nearly all tokens are deactivated in the final layer across five datasets, especially CIFAR-100, SVHN, and K400, where the activation rates are exactly 0%. This indicates that on these datasets, we can directly drop the original MLP block in Layer11 without hurting the performance, which further reduces about 4.7M  [^1] There are two linear layers with weights of size $768\times(4\times 768)$ in a MLP block, results in $2\times 768\times(4\times 768)\approx 4.7$M parameters, 26.6% of the ViT-B total parameters.
Figure 6: Token activation rate in different layers. We visualize the token activation rates from 12 layers in ViT-B/16. “Overall” denotes the mean activation rate in the whole model, which arrives at around 50% when $r$ is set to 0.5. “Layer0” and “Layer11” denote the lowest and highest level, respectively. Notably, the activation rate in the last layer is exactly 0% on CIFAR-100, SVHN, and K400 datasets..
@@
Visualization of activated tokens.
In Figure 7, we visualize two representative samples from K400. We can observe that the model tends to deactivate those tokens that are less informative e.g. tokens of the sky in (a) and tokens of grass in (b). In higher layers, such as layer7 and layer10, only those tokens from the primary objects are activated. This further proves the the existence of token redundancy problems in ViT and provides validation for the rationality behind our approach. Additional visualizations are available in the supplementary material.

### v2

@@ Abstract
Existing parameter-efficient fine-tuning (PEFT) methods have achieved significant success on vision transformers (ViTs) adaptation by improving parameter efficiency. However, the exploration of enhancing inference efficiency during adaptation remains underexplored. This limits the broader application of pre-trained ViT models, especially when the model is computationally extensive. In this paper, we propose Dynamic Tuning (DyT), a novel approach to improve both parameter and inference efficiency for ViT adaptation. Specifically, besides using the lightweight adapter modules, we propose a token dispatcher to distinguish informative tokens from less important ones, allowing the latter to dynamically skip the original block, thereby reducing the redundant computation during inference. Additionally, we explore multiple design variants to find the best practice of DyT. Finally, inspired by the mixture-of-experts (MoE) mechanism, we introduce an enhanced adapter to further boost the adaptation performance. We validate DyT across various tasks, including image/video recognition and semantic segmentation. For instance, DyT achieves superior performance compared to existing PEFT methods while evoking only $71\%$ of their FLOPs on the VTAB-1K benchmark.
@@ 1 Introduction
With the remarkable success of Vision Transformers (ViTs) dosovitskiy2020image ; liu2021swin ; han2022survey , fine-tuing pre-trained ViT on other data domains zhang2023div or task applications zhai2019large ; jia2022visual ; pan2022st ; xie2021segformer is becoming a common practice. However, as model sizes increase zhai2022scaling ; dehghani2023scaling ; liu2022swin , the associated adaptation costs become prohibitive due to the burden of fine-tuning and inference on the target task. Parameter-efficient fine-tuning (PEFT) methods (e.g. AdaptFormer chen2022adaptformer , LoRA hu2021lora , and VPT jia2022visual ), are proposed to tackle the tuning problem by reducing tunable model parameters. They usually update a small amount of parameters while keeping the original model fixed, which reduces learnable parameters effectively while maintaining fine-tuning accuracy.
@@
Figure 1: FLOPs and Accuracy of ViT-B/16 dosovitskiy2020image on VTAB-1K zhai2019large . “Full tuning” denotes that all parameters are fine-tuned. AdaptFormer chen2022adaptformer , LoRA hu2021lora and VPT jia2022visual are typical PEFT methods.
Dynamic networks yin2021adavit ; rao2021dynamicvit ; liang2022not demonstrate that token pruning helps to reduce model inference costs. However, these designs primarily focus on pre-training from scratch or full tuning on the same dataset, without considering data domain transfer. Furthermore, the token pruning operation constraints these methods on visual recognition scenarios, limiting their further applications e.g. dense prediction zhou2017scene . Our motivation thus arises from how to develop dynamic modules together with PEFT methods to enhance both parameter and inference efficiency during ViT adaptation, which should also suit a wide range of visual tasks.
@@
Comprehensive evaluations for DyT are conducted from multiple perspectives. In the image domain, as shown in Figure 1(b). DyT surpasses existing PEFT methods while consuming only $71\%$ of the ViT-B FLOPs on the VTAB-1K benchmark zhai2019large . When visual tokens are scaled up from images to videos, our DyT shows superior generalization ability on action recognition benchmarks, e.g. K400 carreira2017quo and SSV2 goyal2017something , with a reduction of 37GFLOPs. In the scenario where labels are scaled up from recognition to segmentation, our DyT even outperforms full tuning on ADE20K zhou2017scene with 21GFLOPs reduction. These experimental results indicate that the proposed DyT is efficient in both parameter and inference across various visual tasks and directs a promising field for efficient model adaptation.
@@ 2 Related Works
Parameter-efficient fine-tuning (PEFT) is designed to adapt a pre-trained model to downstream tasks by only tuning a small part of the parameters. Existing PEFT methods can broadly be categorized into three groups: adapter-based methods, re-parametrization-based methods, and prompt-based methods. Adapter-based methods chen2022adaptformer ; houlsby2019parameter ; pfeiffer2020adapterfusion ; jie2022convolutional ; chen2022vision insert some tiny modules into the original model and only these inserted modules are updated during fine-tuning. Re-parametrization approaches zaken2021bitfit ; hu2021lora ; lian2022scaling ; edalati2022krona ; dettmers2023qlora usually cooperate with reparameterization techniques, directly modifying specific parameters in the pre-trained model. Prompt-based methods jia2022visual ; bahng2022exploring ; zhou2022learning ; zhou2022conditional ; ju2022prompting ; cao2023domain_quater ; cao2023domain involve appending a small number of learnable tokens to input sequences, thereby adapting intermediate features in frozen layers. Please refer to yu2023visual ; lialin2023scaling to find more comprehensive reviews.
@@
Dynamic neural networkshan2021dynamic ; pu2023adaptive ; zheng2023dynamic can dynamically adjust their architectures based on the input data, which enables them to control the computational redundancy based on input data. Existing methods have explored dynamic layer depth teerapittayanon2016branchynet ; bolukbasi2017adaptive ; yang2020resolution ; han2022learning ; han2023dynamic ; yue2024dynamic , dynamic channel width herrmann2020channel ; li2021dynamic ; han2023latency ; yang2021condensenet and dynamic routing li2020learning in convolutional neural networks. When it comes to the era of vision transformer, many works NeurIPS2020_7866 ; wang2021not ; song2021dynamic ; rao2021dynamicvit ; liang2022not ; wang2021adafocus ; meng2022adavit ; Ni2024AdaNAT ; Ni2024ENAT ; pu2024efficient attempt to improve the inference efficiency by reducing the token redundancy. For instance, Liang et al. liang2022not identify and consolidate inattentive tokens into only one token in each transformer layer. Rao et al. rao2021dynamicvit progressively discard less informative tokens between layers. Wang et al. wang2021not adjust the number of tokens to represent images. A more detailed literature review can be found in han2021dynamic ; wang2023computation . Zhou et al. zhou2024dynamic propose assessing token significance with a ReLU function to effectively tackle point cloud analysis. Although these approaches have shown significant success in vision tasks, they require training a model from scratch or fine-tuning all parameters on the same dataset used in pre-training, making them unsuitable for efficient ViT adaptation.
@@ 3.1 Preliminary
The adapter architectures have been widely explored in previous PEFT methods chen2022adaptformer ; jie2022convolutional ; houlsby2019parameter . An adapter is generally designed as a bottleneck module, which consists of a project layer $\mathbf{W}_{down}\in\mathbb{R}^{C\times d}$ to squeeze the channel dimension and another project layer $\mathbf{W}_{up}\in\mathbb{R}^{d\times C}$ for dimension restoration. Here, $d$ denotes the bottleneck dimension and $d\ll C$ ensures the efficiency of the adapter. Given an input feature $\mathbf{X}$, the adapter operation is formulated as:
@@ 3.2 Dynamic Tuning
In Vision Transformers (ViTs), the computational burden primarily resides in the transformer layers. Specifically, the operations in $\operatorname{Attn}$ and $\operatorname{MLP}$ account for approximately $35.8\%$ and $63.5\%$ of total FLOPs in ViT-B/16 dosovitskiy2020image . Previous works rao2021dynamicvit ; liang2022not have revealed that there exists a token-redundancy issue in ViTs and found that some tokens can be discarded without sacrificing the performance. Inspired by this, we propose an efficient ViT adaptation approach, named Dynamic Tuning (DyT), which not only maintains parameter efficiency during fine-tuning but also reduces redundant computation during inference. The core idea is dynamically selecting tokens for processing within transformer blocks. Given the input tokens $\mathbf{X}$ of a block, DyT can be formulated as:
@@
Figure 2: Overview of Dynamic Tuning. (a) In the fine-tuning stage, we adopt Gumbel Noise to enable end-to-end training. (b)In the inference stage, $\operatorname{TD}$ selects $K$ activated tokens $\mathbf{X}_{s}$ from $\mathbf{X}$ based on the mask $\mathbf{M}$, which saves the computations on those deactivated tokens in $\operatorname{Block}$. $\operatorname{Block}$ can represent a $\operatorname{Attn}$ block, a $\operatorname{MLP}$ block, or an entire transformer layer.
@@
However, directly using threshold makes $\mathbf{M}$ a discrete decision, resulting in a non-differentiable problem during fine-tuning. To address this, we incorporate the Gumbel Noise herrmann2020channel into sigmoid to replace the original sigmoid function during fine-tuning. It can be formulated as:
@@
where $\mathbf{G}_{1}$, $\mathbf{G}_{2}\sim\operatorname{Gumbel}(0,1)$. $\tau$ represents the temperature and is set to 5.0 by default. Further details of Gumbel-Sigmoid are provided in the Appendix A.7. Subsequently, we can obtain $\mathbf{\overline{M}}$ using the same operation in Equation.3. The Gumbel Noise makes the sampling of $\mathbf{\overline{M}}$ stochastic during training and we adopt $\mathbf{\overline{S}}$ as a differentiable approximation of $\mathbf{\overline{M}}$. Both of these help $\operatorname{TD}$ to be trained in an end-to-end manner. The calculation of forward and backward propagations during training can be formulated as:
@@
Figure 3: Model variants. For brevity, we omit the LayerNorm ba2016layer in $\operatorname{Attn}$ and $\operatorname{MLP}$ blocks. “DyT” denotes the dynamic tuning presented in Figure 2.
@@ 3.4 MoE-Adapter
In DyT, the adapter is responsible for processing all tokens, requiring it to have enough capability, particularly when the downstream tasks e.g. semantic segmentation are challenging to adapt. To tackle this problem, we propose a MoE-adapter, inspired by mixture-of-experts shazeer2017outrageously ; yang2019condconv . It effectively enhances the capability of the adapter with introducing negligible computation cost.
@@ 3.5 Loss Functions
For an image $I$, we calculate the cross-entropy loss $\mathcal{L}_{cls}=-\mathbf{y}\operatorname{log}(\mathbf{\hat{y}})$ between the predicted probability $\mathbf{\hat{y}}$ and its corresponding one-hot label $\mathbf{y}$, to supervise the learning of classification. To control the average activation rate $r$ in the entire model, we add a loss term to constrain the number of activated tokens in the dynamic tuning, which can be formulated as: $\mathcal{L}_{rate}=(\frac{1}{L\times N}\sum_{l=1}^{L}\sum_{n=1}^{N}\mathbf{M}_{n}^{l}-r)^{2},$ where $L$ denotes the number of layers in ViT. $\mathbf{M}_{n}^{l}$ represents the mask generated in $\operatorname{TD}$ from the $l$-th layer. Additionally, we employ a loss $\mathcal{L}_{cls}^{\prime}=-\mathbf{y}\operatorname{log}(\mathbf{y}^{\prime})$ to supervise the adaptation of the complete model, where $\mathbf{y}^{\prime}$ is the output probability without employing the token dispatcher. Thus, this complete model can also act as a teacher to enhance the dynamic tuning during adaption by a distillation loss $\mathcal{L}_{distill}=\operatorname{KL}(y^{\prime},y)$, where $\operatorname{KL}$ represents the Kullback-Leibler divergence loss. Therefore, the overall loss function is defined as $\mathcal{L}=\mathcal{L}_{cls}+\mathcal{L}_{cls}^{\prime}+\mathcal{L}_{distill}+\alpha\mathcal{L}_{rate}$, where $\alpha$ serves as the weight of the activation rate loss and is set to 2.0 by default. Note that, DyT can also achieve competitive performance without the distillation loss (see Appendix A.6).
@@ 4.1 Experiment Settings
[…]
@@
We conduct all experiments based on the ViT-B/16 dosovitskiy2020image , which is pre-trained on ImageNet21K deng2009imagenet with full supervision, unless otherwise specified. Results based on ViT-L are presented in the Appendix A.10. The bottleneck dimension $d$ is set to 64 by default. We adopt the same training schedule as chen2022adaptformer . Detailed hyperparameters for each experiment can be found in the Appendix A.8. The default setting in experiments is marked in color.
@@ 4.2 Analysis
In Table 1, we compare the performance of four model variants across both image and video datasets. We set the activation rate $r$ in “Attention Dispatch” and “MLP Dispatch” variants to 0.5, and to 0.7 for “Attention-MLP Dispatch” and “Layer Dispatch” variants, and train each respectively. This results in similar average FLOPs for four variants. We observe that the default setting “MLP Dispatch” achieves superior performance across five datasets while maintaining the lowest computational cost. Although “Attention-MLP Dispatch” and “Layer Dispatch” also exhibit good performance on K400, the former incurs double additional parameters while the latter lacks generalization capability on other datasets. The comparison between “MLP Dispatch” and other variants proves that only skipping tokens in MLP blocks is a better design. More investigations on model variants can be found in our Appendix A.3.
@@
We observe that the performance on image datasets drops when we increase the expert number in MoE-adapter. This phenomenon can be attributed to the simplicity of image datasets and the model does not require too many parameters to adapt. In contrast, for video datasets, such as K400 and SSv2, the best accuracy is achieved 4 experts. The reason behind this is that the domain gap between the pre-training dataset and video datasets is large and the model needs sufficient adapter capacity to learn the adaptation. This proves that we can introduce the MoE-adapter when the target dataset is difficult to adapt.
@@
In Figure 5, we visualize the token activation rates across different layers in ViT-B/16 dosovitskiy2020image . We observe that the model tends to activate more tokens in lower layers while deactivating tokens in higher layers. This phenomenon can be attributed to that the model tends to take more general knowledge from the lower layers of the pre-trained model and learn more task-specific knowledge in higher levels. Additionally, the activation results vary across different datasets. For instance, SSv2 demonstrates increased token activation rate in Layer 5 and Layer 6 compared to other datasets, whereas SVHN experiences substantial token deactivation in Layer 6, 7, and 8. This discrepancy arises from that the model needs different knowledge from the pre-trained weights to address dataset-specific challenges.
It is noteworthy that nearly all tokens are deactivated in the final layer across five datasets, especially CIFAR-100, SVHN, and K400, where the activation rates are exactly 0%. This indicates that on these datasets, we can directly drop the original MLP block in Layer11 without hurting the performance, which further reduces about 4.7M ***There are two linear layers with weights of size $768\times(4\times 768)$ in a MLP block, results in $2\times 768\times(4\times 768)\approx 4.7$M parameters, 5.4% of the ViT-B total parameters.
Figure 5: Token activation rate in different layers. We visualize the token activation rates in ViT-B/16. “Overall” denotes the mean activation rate in the whole model, which arrives at around 50% when $r$ is set to 0.5. “Layer0” and “Layer11” denote the lowest and highest level, respectively. Notably, the activation rate in the last layer is exactly 0% on CIFAR-100, SVHN, and K400 datasets.
Visualization of activated tokens.
In Figure 6, we visualize two representative samples from K400. We can observe that the model tends to deactivate those tokens that are less informative e.g. tokens of the sky in (a) and tokens of grass in (b). In higher layers, such as layer7 and layer10, only those tokens from the primary objects are activated. This further proves the the existence of token redundancy problems in ViT and provides validation for the rationality behind our approach. Additional visualizations are available in Appendix A.11.